# HSSBench: Benchmarking Humanities and Social Sciences Ability for Multimodal Large Language Models

### v1

@@ 5 Related Work
In recent years, significant progress has been made in the development of multimodal benchmarks and methodologies. Numerous datasets have emerged that assess models from various perspectives, which can be broadly categorized into Generation Benchmarks liu2024mmbench ; liu2023visualinstructiontuning ; meng2024vgavisionguiassistant ; han2024instinctivebiasspuriousimages ; qian2024easyfoolmultimodalllms ; xu2023lvlmehubcomprehensiveevaluationbenchmark ; yin2023lammlanguageassistedmultimodalinstructiontuning ; zeng2023matterstraininggpt4stylelanguage ; wu2024qbenchbenchmarkgeneralpurposefoundation ; luo-etal-2024-codis , Reasoning Benchmarks wang2024measuringmultimodalmathematicalreasoning ; lu2024mathvistaevaluatingmathematicalreasoning ; li2023scigraphqalargescalesyntheticmultiturn ; liang2024scemqascientificcollegeentrance ; song2025csbench , and Application Benchmarks fan2022minedojobuildingopenendedembodied ; rawles2023androidwildlargescaledataset ; chen2023endtoendembodieddecisionmaking ; sermanet2023robovqamultimodallonghorizonreasoning ; chen2024egoplanbenchbenchmarkingmultimodallarge ; koh2024visualwebarenaevaluatingmultimodalagents ; you2024ferretuigroundedmobileui . Generation benchmarks cover a wide range of tasks, with the core objective of comprehensively evaluating the multi-dimensional performance of MLLMs. They not only include reasoning challenges but also address areas related to humanities and social sciences. In contrast, reasoning benchmarks are designed primarily around mathematical and scientific problems, seeking to rigorously assess the capabilities of MLLM in logical and analytical reasoning. Among the various evaluation formats, multiple-choice datasets are the most prevalent, owing to their simplicity in evaluation and ease of comparison zeng2023matterstraininggpt4stylelanguage .
@@
The proliferation of these data sets has significantly accelerated research progress, serving both as training resources and as tools for assessing the multifaceted competencies of MLLMs. Advances in data set construction are particularly exciting, ranging from human-annotated high-quality datasets romero2024cvqaculturallydiversemultilingualvisual to those generated through pipelines built in LLM chandrasegaran2024hourvideo1hourvideolanguageunderstanding . Despite these achievements, several critical limitations remain. 1) Most multimodal datasets cover a wide range of categories, but suffer from issues such as limited data sources, relatively simple questions, and insufficient image information. 2) Many large-scale datasets are collected through web crawling without thorough manual annotation and verification, which can introduce biases in evaluation results. 3) Current reasoning datasets predominantly focus on STEM tasks, relying heavily on scientific and mathematical data to assess reasoning capabilities. In contrast, the benchmark we propose is created through interactions among expert agents, targeting the HSS domain. In contrast, the benchmark we propose focuses on the HSS domain, emphasizing tasks that analyze and understand the abstract concepts embedded in images. This approach fills a notable gap in the field and lays the foundation for advancing model performance in this underexplored area.
@@ C.1 Details of Results
Open-source LLM (Scale <<< 10B)
@@
Open-source LLM (Scale >>> 10B)
@@
Table 4: Scores (%) of MLLMs on HSSBench (EN-\@slowromancapii@).
@@
C.2 Details of Contrast Experiment
@@
C.3 Results in different languages
@@
Model Geography Economics Culture Social Sciences History Art All
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale <<< 10B)
Open-source LLM (Scale >>> 10B)
Closed-source LLM
Table 7: Scores (%) of MLLMs on HSSBench (ZH-\@slowromancapi@).
@@
Model Geography Economics Culture Social Sciences History Art All
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale <<< 10B)
Open-source LLM (Scale >>> 10B)
Closed-source LLM
Table 8: Scores (%) of MLLMs on HSSBench (ZH-\@slowromancapii@).
Model Arabic Chinese English French Russian Spanish
InternVL3-8B 35.20 49.37 42.38 39.37 38.70 39.21
Qwen2.5-VL-7B-Instruct 33.74 40.21 38.37 34.08 35.55 34.83
Qwen2.5-VL-32B-Instruct 46.94 54.79 50.24 47.57 48.51 50.24
Qwen2.5-VL-72B-Instruct 49.72 55.41 51.28 51.28 48.27 48.83
QVQ-72B-Preview 33.98 36.89 39.80 37.86 40.77 38.83
GPT-4.1-mini 41.33 44.89 41.67 41.78 41.44 46.89
Average 40.82 46.93 43.96 41.99 42.87 43.47
C.4 Prompt for Model Inference
@@
C.5 Evaluation for model’s output
@@ Appendix D More Cases
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have detailed in the abstract and introduction that HSS-related issues still pose significant challenges for large models. Our contribution lies in proposing a large-scale, high-quality benchmark to evaluate the performance of large models on HSS tasks.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section 4, we discuss two limitations of our work: 1) Since there is no fixed set of classification tasks in the HSS field, our dataset does not cover all possible tasks within HSS; 2) Some data in certain HSS subfields have spatiotemporal relevance, so we can only ensure that the questions in our dataset are correct at the present time.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work is focused on datasets and does not involve theoretical assumptions or proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our experimental setup prompts are described in detail in Appendix C.4 and  C.5. We also verify the consistency of evaluation across different models; that is, similar results can be achieved using other open-source models as with our closed-source model tests. Our work can be reproduced on our dataset from both the testing and evaluation perspectives.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have open-sourced our dataset and implementation code on both GitHub and Hugging Face.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 3.1, we provide a detailed description of our experimental design. Our experimental setup prompts are described in detail in Appendix C.4 and  C.5.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In the Appendix C.5, we discuss the consistency among human evaluation, GPT evaluation, and Qwen evaluation. The consistency is within an acceptable range, demonstrating that the choice of evaluation method does not have a significant statistical impact on the final results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: Our benchmark work is not related to computing resources; as long as it can run, that’s sufficient.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work complies with ethical standards, and all experts involved in data production are members of our team. Moreover, the images we use are either drawn by our experts or sourced from open-source data or websites. Our research does not have any harmful social impact.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Throughout this paper, we emphasize the positive social impact of our work. Our evaluation dataset can help the AI community assess the performance of large models on HSS tasks, while also highlighting the current insufficient attention given to HSS tasks. We believe our work can actively advance the development of MLLMs.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: The construction of our dataset involves the use of multiple agents for web data crawling. However, we have implemented numerous restrictions during the crawling process, and image and text information can only be obtained from websites that we have selected and approved. We assure you that there are no risks associated with the use of our networked multi-agent system.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We obtained data from https://www.unesco.org/zh/open-access/cc-nc-sa and comply with the CC BY-NC-SA license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our dataset contains 13,152 image-text pairs, and each entry supports evaluation in the six official languages of the United Nations
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our dataset work does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our dataset work does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@ 4 Related Work
In recent years, significant progress has been made in the development of multimodal benchmarks and methodologies. Numerous datasets have emerged that assess models from various perspectives, which can be broadly categorized into Generation Benchmarks Liu et al. (2024b; 2023); Meng et al. (2024); Han et al. (2024); Qian et al. (2024); Xu et al. (2023); Yin et al. (2023); Zeng et al. (2023); Wu et al. (2024a); Luo et al. (2024); Feng et al. (2025); Wang et al. (2025e), Reasoning Benchmarks Wang et al. (2024a); Lu et al. (2024); Li and Tajbakhsh (2023); Liang et al. (2024); Song et al. (2025), and Application Benchmarks Fan et al. (2022); Rawles et al. (2023); Chen et al. (2023); Sermanet et al. (2023); Chen et al. (2024a); Koh et al. (2024); You et al. (2024); Gao et al. (2025); Kang et al. (2026); Zheng et al. (2026); He et al. (2026). Generation benchmarks cover a wide range of tasks, often including not only reasoning challenges but also content related to the humanities and social sciences. Their main objective is to provide a comprehensive assessment of the performance of MLLMs in various dimensions. In contrast, reasoning benchmarks are designed primarily around mathematical and scientific problems, seeking to rigorously assess the capabilities of MLLM in logical and analytical reasoning. Among the various evaluation formats, multiple-choice datasets are the most prevalent, owing to their simplicity in evaluation and ease of comparison Zeng et al. (2023).
@@
The proliferation of these data sets has significantly accelerated research progress, serving both as training resources and as tools for assessing the multifaceted competencies of MLLMs Wang et al. (2025f); Liu et al. (2025b); Luo et al. (2026); Guan et al. (2026); Fu et al. (2026); Wu et al. (2025; 2024b); Feng et al. (2026); Li et al. (2026; 2025c; 2025d); Hu et al. ; Zhang et al. (2025c; b; a); Ding et al. (2025); Li et al. (2025b); Wang et al. (2026); Ma et al. (2026); Liu et al. (2025a); Bi et al. (2026); Wang et al. (2025d; c; b; a); Liang et al. (2025); Ma et al. (2025a; b); Li et al. (2025a); Yu et al. (2025); Feng et al. (2024; 2023). Advances in data set construction are particularly exciting, ranging from human-annotated high-quality datasets Romero et al. (2024) to those generated through pipelines built in LLM Chandrasegaran et al. (2024). Despite these achievements, several critical limitations remain. 1) Most multimodal datasets cover a wide range of categories, but suffer from issues such as limited data sources, relatively simple questions, and insufficient image information. 2) Many large-scale datasets are collected through web crawling without thorough manual annotation and verification, which can introduce biases in evaluation results. 3) Current reasoning datasets predominantly focus on STEM tasks, relying heavily on scientific and mathematical data to assess reasoning capabilities. In contrast, the benchmark we propose is created through interactions among expert agents, targeting the HSS domain. In contrast, the benchmark we propose focuses on the HSS domain, emphasizing tasks that analyze and understand the abstract concepts embedded in images. This approach fills a notable gap in the field and lays the foundation for advancing model performance in this underexplored area.
@@ A.1 Annotation Platform
A.2 Information about Experts
Name Languages Professional Background
Expert 1 Chinese, English Computer Science, Art, Economy
Expert 2 Chinese, Japanese, English Computer Science, Culture
Expert 3 English, Chinese Computer Science, Geography
Expert 4 Chinese, English Economy, History
Expert 5 Chinese, English, Japanese Language and Linguistics, History, Culture
Expert 6 Chinese, English, Russian Language and Linguistics, Social Science
Expert 7 Chinese, English, Arabic Language and Linguistics, Social Science
Expert 8 Chinese, English, Arabic Economy, Social Science
Expert 9 Chinese, English Geography
Expert 10 Chinese, English, French Economy, Art
Expert 11 English, Chinese, Japanese Art, Culture
Expert 12 Chinese, English Computer Science, Geography
Expert 13 Chinese, English Geography, History
Expert 14 Chinese, English, Spanish Computer Science, Geography
Table 2: Information about the experts involved in dataset construction.
The team of experts who contributed to the creation of HSSBench comes from a variety of cultural and linguistic backgrounds, as well as interdisciplinary academic fields spanning the humanities and social sciences. This diversity is reflected in their linguistic fluency, professional expertise, and international academic experiences. Table 2 summarizes the profiles of the experts involved in the dataset construction.
Many of the experts have international academic experiences and interdisciplinary training across various humanities and social science domains. While individual cultural perspectives may naturally influence the emphasis or framing of certain questions, we consider this diversity a strength rather than a limitation. It enriches the dataset by incorporating multiple viewpoints and insights, which is essential for a benchmark designed to be multilingual and cross-cultural.
To mitigate potential cultural bias, we carefully curated the source materials provided to the experts during the initial stage of dataset construction. These reference materials originate from multiple countries and languages, ensuring that the constructed questions and answers reflect widely accepted, global knowledge rather than culturally subjective viewpoints. Furthermore, during the data validation and filtering stages, we rigorously excluded content that could be culturally biased or inconsistent with universal values.
In summary, our approach balances the preservation of valuable cultural diversity with the need to maintain fairness and universality in the dataset. This careful design ensures that HSSBench serves as a robust and inclusive benchmark for evaluating multilingual large language models (MLLMs) across different languages and cultural contexts.
@@ B.3 Most-Frequent Words in the Questions
B.4 Composition and Quality Comparison Between Human and Automated Agents
We provide a detailed breakdown of the data sources across the six domains in HSSBench, as shown in Table 5.
Category Human Expert Automated Agent Total Samples
Art 984 1,204 2,188
Geography 2,671 891 3,562
History 1,946 550 2,496
Economy 974 459 1,433
Social Science 487 947 1,434
Culture 652 1,387 2,039
Table 5: Distribution of samples constructed by human experts and automated agents across different domains in HSSBench.
To further validate the quality of data generated by both human experts and automated agents, we evaluated the performance of GPT4.1-mini on the respective subsets. The accuracy results are summarized in Table 6.
Category Human Expert Acc. Automated Agent Acc.
Art 42.32 45.43
Geography 48.82 46.52
History 45.59 49.12
Economy 58.87 57.67
Social Science 43.15 46.95
Culture 49.11 47.41
Table 6: Model accuracy (%) of GPT4.1-mini on human-expert-generated and automated-agent-generated subsets across different domains.
The comparable accuracy results indicate that data generated by automated agents, after undergoing multi-round expert verification, maintain a quality level consistent with that of human expert contributions. These clarifications and statistics have been incorporated into the revised manuscript to enhance transparency regarding dataset construction and quality assurance.
@@ C.1 Details of Results
Open-source LLM (Scale << 10B)
@@
Open-source LLM (Scale >> 10B)
@@
Table 7: Scores (%) of MLLMs on HSSBench (EN-II).
@@
C.2 Results in different languages
Model Geography Economics Culture Social Sciences History Art All
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale << 10B)
Open-source LLM (Scale >> 10B)
Closed-source LLM
Table 8: Scores (%) of MLLMs on HSSBench (ZH-I).
Model Geography Economics Culture Social Sciences History Art All
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale << 10B)
Open-source LLM (Scale >> 10B)
Closed-source LLM
Table 9: Scores (%) of MLLMs on HSSBench (ZH-II).
C.3 Multiple-Choice Confounding Option Experiment Details
Model Arabic Chinese English French Russian Spanish
InternVL3-8B 35.20 49.37 42.38 39.37 38.70 39.21
Qwen2.5-VL-7B-Instruct 33.74 40.21 38.37 34.08 35.55 34.83
Qwen2.5-VL-32B-Instruct 46.94 54.79 50.24 47.57 48.51 50.24
Qwen2.5-VL-72B-Instruct 49.72 55.41 51.28 51.28 48.27 48.83
QVQ-72B-Preview 33.98 36.89 39.80 37.86 40.77 38.83
GPT-4.1-mini 41.33 44.89 41.67 41.78 41.44 46.89
Average 40.82 46.93 43.96 41.99 42.87 43.47
@@ C.4 Visual Information Extraction Experiment Details
C.5 Comparison with HSS-Related Benchmarks
We acknowledge that some existing benchmarks, such as MME, include HSS-related test data, particularly within the Art domain. To better understand the relationship between these datasets and our proposed HSSBench, we conducted a detailed comparative analysis focusing on the overlapping Art category. The evaluation was performed under identical prompt settings, including both Direct and Chain-of-Thought (CoT) prompting. The combined results are presented in Table 13.
Model Benchmark (% Accuracy) Prompting
CMMMU MME MMMU HSSBench Direct CoT
Qwen2.5-VL-3B-Instruct 47.73 77.32 57.76 29.01 47.73 48.86
Qwen2.5-VL-7B-Instruct 43.18 71.21 61.21 37.88 43.18 46.59
InternVL3-8B 65.91 85.28 68.10 42.14 65.91 55.68
llava-onevision-qwen2-7b 53.41 77.21 49.14 36.20 53.41 52.27
Table 13: Performance comparison on the Art category across HSS-related benchmarks under Direct and CoT prompting.
As shown, the relative performance trends across these benchmarks are broadly consistent, indicating that the challenges within the Art domain are similar across datasets. Notably, HSSBench presents a more challenging evaluation, reflected by generally lower accuracy scores, which underscores its value in pushing forward research on Humanities and Social Sciences tasks.
C.6 Comparison with STEM Benchmarks
While our primary focus is on Humanities and Social Sciences (HSS), we also recognize the importance of situating HSSBench within the broader landscape of STEM benchmarks. To this end, we provide a comparative overview of model performance on several representative STEM benchmarks alongside HSSBench under the Chain-of-Thought (CoT) prompting setting. The results are summarized in Table 14.
Model MMLU Pro GPQA Diamond SWE-bench Verified MATH- 500 AIME 2024 LiveCode Bench OpenCompass- Reasoning HSSBench
Qwen2.5-VL-72B 71.2 50.2 51.87
GPT-4.1 mini 65.00 23.60 49.60 46.0 48.03
GPT-4o 79.80 66.90 35.80 54.8 46.88
GPT-4.1 81.80 66.30 54.6 92.80 48.10 40.50 54.0 42.66
InternVL3-8B 41.4 41.42
GPT-4.1 nano 50.30 29.40 34.2 35.83
Janus-Pro-7B 19.1 31.66
Table 14: Model performance comparison on STEM benchmarks and HSSBench (CoT prompting).
Notes: All STEM benchmark and OpenCompass results are sourced from official publications or model release notes. A dash (—) indicates that the model did not publicly report results for that benchmark. Minor discrepancies may exist due to testing variability but do not affect the overall trend.
This comparison reveals that while models generally achieve higher accuracy on STEM benchmarks (e.g., MMLU Pro, MATH-500), their performance on HSSBench is comparatively lower. This gap highlights the unique challenges posed by HSS tasks and the necessity of dedicated benchmarks like HSSBench to drive progress in this domain.
C.7 Retrieval-Augmented Generation Evaluation
We conducted additional experiments integrating Retrieval-Augmented Generation (RAG) with several smaller MLLMs. The retrieval database was constructed from a general knowledge corpus comprising Wikipedia and publicly available documents related to HSS. Model performance was evaluated under two prompting strategies: direct prompting and CoT prompting.
Table 15: Performance of models under direct prompting.
@@ C.7 Retrieval-Augmented Generation Evaluation
Model Geography Art Culture Social Science History Economy Overall
Without RAG
Qwen2.5-VL-7B-Instruct 60.09 28.05 27.19 53.50 54.51 40.09 43.89
Qwen2.5-VL-3B-Instruct 55.40 33.48 44.24 54.50 50.82 35.59 45.56
InternVL3-8B 54.93 37.56 43.78 62.00 51.64 44.14 48.82
llava-onevision-qwen2-7b 46.01 30.77 34.10 46.50 38.93 29.28 37.43
With RAG
Qwen2.5-VL-7B-Instruct 47.42 29.86 29.95 51.00 45.08 38.74 40.24
Qwen2.5-VL-3B-Instruct 49.77 33.03 35.48 48.50 43.44 33.78 40.55
InternVL3-8B 53.52 35.75 35.94 59.50 49.59 37.39 45.10
llava-onevision-qwen2-7b 43.19 32.13 26.73 41.00 32.38 21.17 32.57
Table 16: Performance of models under COT prompting.
@@
Model Geography Art Culture Social Science History Economy Overall
Without RAG
Qwen2.5-VL-7B-Instruct 55.40 25.79 27.65 51.00 47.54 39.64 41.08
Qwen2.5-VL-3B-Instruct 51.17 30.32 35.48 49.00 46.31 32.43 40.70
InternVL3-8B 53.52 34.39 40.55 57.00 51.64 38.74 45.86
llava-onevision-qwen2-7b 43.66 31.22 28.57 43.00 33.61 29.73 34.78
With RAG
Qwen2.5-VL-7B-Instruct 47.89 27.15 27.19 46.50 42.62 36.49 37.89
Qwen2.5-VL-3B-Instruct 41.78 31.22 29.95 44.00 40.57 36.04 37.21
InternVL3-8B 57.75 30.77 34.10 55.50 50.41 39.19 44.50
llava-onevision-qwen2-7b 42.25 29.86 28.11 42.50 31.56 23.42 32.73
The results indicate that, although RAG occasionally yields modest improvements in specific domains or models, it does not consistently outperform direct prompting without retrieval augmentation. This suggests that augmenting MLLMs with a general retrieval corpus and straightforward prompting strategies may be insufficient to fully exploit the complex and nuanced knowledge required for HSS tasks.
These findings highlight the challenges inherent in applying RAG to Humanities and Social Sciences benchmarks such as HSSBench. We hypothesize that more specialized, domain-specific retrieval corpora, combined with advanced retrieval and integration techniques, are necessary to unlock the full potential of RAG in this context.
C.8 Prompt for Model Inference
@@
C.9 Evaluation for model’s output